OpenStudy (anonymous):

What is the slope-intercept form of this equation? 8x + 6y = 12

OpenStudy (igreen):

8x + 6y = 12 Slope-Intercept Form is y = mx + b. Just rearrange the terms. Subtract 8x to both sides: 6y = -8x + 12 Divide 6 to all terms: \(y = -\dfrac{8}{6}x + \dfrac{12}{6}\) Just simplify the 12 / 6 at the back. What's 12 / 6? @RainbowBubbles01

OpenStudy (igreen):

@RainbowBubbles01

OpenStudy (anonymous):

its 12 over

OpenStudy (anonymous):

6

OpenStudy (anonymous):

@iGreen